# Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes

### v3

@@ 21 W32 LPT coolant bleed
7.1 Comparison to Baselines
To analyze our methodology, we compare its performance to that of [39] Chehade et al. (2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [40] Li et al. (2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection. Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles. From this figure, our methodology becomes more accurate as the RUL decreases. Furthermore, our methodology is more accurate than both when making longer predictions. For test units whose RUL is at most 20, our accuracy is between that of [39] and [40]. Only at RUL at most 40, does our methodology perform worse than both. To further analyze our methodology, we refer to Figure 8, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: $[0,25],[26,50],[51,75],[76,100],[101,125],[126,150]$. The results indicate that our methodology achieves a lower median relative error than [39] in all intervals while maintaining comparable variability. However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [39]. This indicates a targeted area for improvement of our methodology. However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection. Furthermore, we were capable of achieving more accurate RUL prediction than both methods from earlier in the systems’ lives despite not having information about the actual failure mode.
@@ 8 Conclusion
Acknowledgments
This effort is supported by NASA under grant number 80NSSC19K1052 as part of the NASA Space Technology Research Institute (STRI) Habitats Optimized for Missions of Exploration (HOME) ‘SmartHab’ Project. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of the National Aeronautics and Space Administration.
@@ Appendix A Log Likelihood
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\times\ \frac{\rho_{k}}{\sqrt{2\pi}}\exp\left(-\frac{1}{2}\left(\rho_{k}y_{i}-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k}\right)^{2}\right)$ (25)
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vspace*{-1em}$
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vspace*{-1em}$
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vspace*{-1em}$

### v4

@@ Appendix A Log Likelihood
$\begin{split}\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\sum_{k=1}^{K}\pi_{k}\frac{\rho_{k}}{\sqrt{2\pi}}&\exp(-\frac{1}{2}(\rho_{k}y_{i}\\ &-\psi_{0,k}-\sum_{p=1}^{P}\boldsymbol{x}_{i,p}^{T}\boldsymbol{\psi}_{p,k})^{2})\end{split}$ (25)
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})=\sum_{i=1}^{N}\ln\left(\sum_{k=1}^{K}\frac{g\left(\boldsymbol{Z}_{i}[k]\right)p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}\right)\vskip-10.00002pt$
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{\frac{p\left(y_{i},\boldsymbol{Z}_{i}[k];\boldsymbol{\Theta}\right)}{g\left(\boldsymbol{Z}_{i}[k]\right)}}\vskip-10.00002pt$
@@
$\ell(\boldsymbol{\Theta};\boldsymbol{Y})\geq\mathbb{E}_{\boldsymbol{g}}\left[\ell_{C}(\boldsymbol{\Theta};\boldsymbol{Y},\mathbb{Z})\right]-\sum_{i=1}^{N}\sum_{k=1}^{K}g\left(\boldsymbol{Z}_{i}[k]\right)\ln{g\left(\boldsymbol{Z}_{i}[k]\right)}\vskip-10.00002pt$